OpenStudy (anonymous):

A mysterious box is delivered to the dinner party you are attending. The label on the box says that the volume of a box is the function f(x) = x3 + 3x2 – 10x – 24. To open the box, you need to identify the correct factors of f(x). Partygoers offer up solutions, and it is your job to find the right ones. Their suggestions are: (x – 1) (x + 2) (x – 3) (x + 4) (x + 6) (x – 12) List the correct factors. Then justify your selections with complete sentences.

OpenStudy (marcon):

(x+2) (x-3) (x+4)

OpenStudy (anonymous):

thanks but how would i justify with complete sentences ?

OpenStudy (marcon):

I'm not sure how many sentences you need, but I would just say: Find the Greatest Common Factor (GCF), then factor from that For example lets take 24 and factor that out 24 ^ 6 4 ^ ^ 3 2

OpenStudy (anonymous):

i would only need three sentences just for each answer you gave me i would need an explanation of how each answer you gave me works. so basically i would just need to show my work

OpenStudy (marcon):

I showed you an example how to show your work, you should be able to do the rest on your own. Just find the GCF and factor thats all there is to it, really

OpenStudy (anonymous):

ok thanks

OpenStudy (marcon):

no problem

OpenStudy (anonymous):

can you help with these questions ? Three partygoers are in the corner of the ballroom having an intense argument. You walk over to settle the debate. They are discussing a function g(x). You take out your notepad and jot down their statements. Professor McCoy: She says that 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. Ms. Guerra: She says that 2 is a zero of g(x) because g(2) = 0. Mr. Romano: He says that 2 is a zero of g(x) because synthetic division with 2 results in a remainder of 0. Correct the reasoning of any inaccurate reasoning by the partygoers in full and complete sentences. Make sure you reference any theorems that support your justifications. Dr. Collier summons you over to his table. He wants to demonstrate the graph of a fourth-degree polynomial function, but the batteries in his graphing calculator have run out of juice. Explain to Dr. Collier how to create a rough sketch of a graph of a fourth-degree polynomial function. Mrs. Collins is at the table with you and states that the fourth-degree graphs she has seen have 4 real zeros. She asks you if it is possible to create a fourth-degree polynomial with only 2 real zeros. Demonstrate how to do this and explain your steps.

OpenStudy (marcon):

Sorry I can't, but I can get someone who might help you out... @jim_thompson5910

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Also im new to open study so how could I get him to help me ?

OpenStudy (marcon):

Just tag some people who are online or people who helped you before. I'm pretty new also, created my account last week lol

OpenStudy (anonymous):

lol i just created mine like 30 minutes ago bro but ok thanks.

OpenStudy (marcon):

yeah sure

OpenStudy (anonymous):

Lincoln and Gemma are looking at the equation the square root of the quantity of 2 times x minus 3 equals square root of x . Lincoln says that the solution is extraneous. Gemma says the solution is non-extraneous. Is Lincoln correct? Is Gemma correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences. can you help with this ?

OpenStudy (marcon):

I did this question in a test before, so I'll just paste it all for u. √2x - 3 = √x Square both sides. 2x - 3 = x Subtract x from both sides and you get x - 3 = 0 Add 3 on each side x = 3 Substitute 3 into the original equation in for x √2(3) - 3 = √3 √3 = √3 Gemma is correct, and Lincoln is wrong. The solution is non-extraneous

OpenStudy (marcon):

Refresh the page if there's a bunch of question marks instead of the equation